OpenStudy (anonymous):

5x^-2y^10 over 2x^-1(-3x^-3y^-1)^-2 very hard for me

OpenStudy (anonymous):

i just dont know i got this answer (45/2)x^-7 y^8

mathslover (mathslover):

are there any options

OpenStudy (anonymous):

no

mathslover (mathslover):

gr8 .. :(

OpenStudy (anonymous):

its super hard

mathslover (mathslover):

not super hard it is super complex

OpenStudy (callisto):

\[\frac{5x^{-2}y^{10}}{2x^{-1}(-3x^{-3}y^{-1})^{-2}}\]\[=\frac{5x^{-2}y^{10}}{2x^{-1}((-3)^{(-2)}x^{(-3)(-2)}y^{(-1)(-2)})}\]\[=\frac{5x^{-2}y^{10}}{2x^{-1}((9)^{(-1)}x^{6}y^{2}}\]\[=\frac{5x^{-2}y^{10}}{2(9)^{(-1)}x^{5}y^{2}}\]\[=\frac{5(9)y^{10-2}}{2x^{5+2}}\]\[=\frac{45y^{8}}{2x^{7}}\]\[=\frac{45}{2}x^{-7}y^{8}\]To be frank, it's not that difficult...

mathslover (mathslover):

gr8 @Callisto

OpenStudy (anonymous):

was i right

mathslover (mathslover):

ya u were right gr8 work @demarco1322

OpenStudy (callisto):

Wait... should you leave the answer in positive indices?

OpenStudy (anonymous):

thanks the both of you

OpenStudy (callisto):

If you need, then the answer is the second last step; if not, the your answer is okay :)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

Welcome!~

OpenStudy (anonymous):

\[\frac{(5x{}^{\wedge}-2y{}^{\wedge}10)}{ (2x{}^{\wedge}-1(-3x{}^{\wedge}-3y{}^{\wedge}-1){}^{\wedge}-2)}=\frac{\frac{5 y^{10}}{x^2}}{ \left(\frac{2}{x}\right)\left(\frac{x^6 y^2}{9}\right)} =\frac{\frac{5 y^{10}}{x^2}}{ \frac{2 x^5 y^2}{9}}=\frac{45 y^8}{2 x^7} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your are welcome.